OpenStudy (anonymous):

what is the intersection point of the graphs of the equations f(x)=4x-1 and g(x)=x+2

OpenStudy (anonymous):

If they intersect that means that they have the same value so: f(x)=g(x) 4x-1 = x+2 3x = 3 x=1 plug that in any of the two functions to get the y coordinate